OpenStudy (anonymous):

simplify by removing factors of 1: (m^2-1)/(m+1)^2

jimthompson5910 (jim_thompson5910):

\[\large \frac{m^2-1}{(m+1)^2}\] \[\large \frac{(m-1)(m+1)}{(m+1)(m+1)}\] \[\large \frac{m-1}{m+1}\] So \[\large \frac{m^2-1}{(m+1)^2}=\frac{m-1}{m+1}\]